OpenStudy (anonymous):

-4(x + 3) - 7 = 4x + 13 i've already gotten to -4x-12-7=4x+13

Nnesha (nnesha):

good now combine like terms which is -12-7

OpenStudy (anonymous):

-4x-19=4x+13

Nnesha (nnesha):

after that subtract 4x both side let me know what you get after that @bree_ramos

OpenStudy (anonymous):

-4x(+4x)-19(-13)=(4x+)4x+13(-13)

OpenStudy (anonymous):

then cancel out the rest of the stuff.

Nnesha (nnesha):

do you understand ?? if not let me know i will try my best to explain :) :) :)

OpenStudy (anonymous):

i was following rosho's way and i got 32. idk if it's right or not tho.

Nnesha (nnesha):

is that final answer ??

OpenStudy (anonymous):

yeah that's what i got.

Nnesha (nnesha):

nope that's not right lets do that again so you got -4x -12 -7 =4x + 13

Nnesha (nnesha):

now combine like terms -12-7 =???

Nnesha (nnesha):

-4x -19 =4x + 13 right ?? did you get this before

OpenStudy (anonymous):

yes.

Nnesha (nnesha):

so now add 4x both side let me know what you get after that

Nnesha (nnesha):



OpenStudy (anonymous):

yes i did that

Nnesha (nnesha):

ok so after that subtract 13 both side =???

OpenStudy (anonymous):

so x= 4 (: ?

Nnesha (nnesha):

yeah :O but you forgot something

OpenStudy (anonymous):

-4!

Nnesha (nnesha):

yep!!! goodi job

OpenStudy (anonymous):

thank you for your (: haha i might need it again in a minute

Nnesha (nnesha):

my pleasure sure i will help you again if i can :(